OpenStudy (anonymous):

Help me Por Favor....Choose the correct slope of the line that passes through the points (1, -3) and (3, -5) .

OpenStudy (sleepyhead314):

Hello and Welcome to OpenStudy! :) \[slope = \frac{ rise }{ run } = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} } \] when given points (x1, y1) and (x2, y2)

OpenStudy (anonymous):

Identify the horizontal asymptote of f(x) = quantity 7 x plus 1 over quantity 2 x minus 9

OpenStudy (sleepyhead314):

@alexandermarrero17 please try posting your own questions in the "Ask a question..." box to the left :)

OpenStudy (camerondoherty):

use the point slope formula: \(\Large\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\) (1, -3) and (3, -5) \(\Large\frac{-5 - (-3) }{3 - 1 }\)

OpenStudy (anonymous):

o ok my bad

OpenStudy (camerondoherty):

since its a double negative you add: \[\frac{ -5+3 }{ 3-1 }\]

OpenStudy (camerondoherty):

Now you simplify that

OpenStudy (anonymous):

Options are A. -1 B.1 C.0 D.1/2 I got A...I'm just tryna see if I'm right

OpenStudy (anonymous):

@camerondoherty

OpenStudy (camerondoherty):

-5 + 3 = -2 3 - 1 = 2 \[\frac{ -2 }{ 2 }=?\]

OpenStudy (anonymous):

so A ?

OpenStudy (anonymous):

Am I right

OpenStudy (camerondoherty):

What does it equal?

OpenStudy (camerondoherty):

Whatever it equals is your answer

OpenStudy (anonymous):

Oh yee...So I am right :)

OpenStudy (anonymous):

Thank you @camerondoherty... Ima need you in the future tbh